# A Semantic-Aware and Multi-Guided Network for Infrared-Visible Image Fusion

### v2

@@ Abstract
Multi-modality image fusion aims at fusing modality-specific (complementarity) and modality-shared (correlation) information from multiple source images. To tackle the problem of the neglect of inter-modal information, high-frequency information loss and limited attention to downstream tasks, this paper focuses on how to model correlation-driven decomposing features and reason high-level graph representation by efficiently extracting complementary features and aggregating multi-guided features. We propose a three-branch encoder-decoder architecture along with corresponding fusion layers as the fusion strategy. The transformer with Multi-Dconv Head Transposed Attention and Local-enhanced Feed Forward network is used to extract shallow features after the depthwise convolution. In the three parallel branches of the encoder, Cross Attention and Invertible Block (CAI) enables to extract local features and preserve high-frequency texture details. Base Feature Extraction Module (BFE) with residual connections can capture long-range dependencies and enhance modality-shared information. Graph Reasoning Module (GR) is introduced to reason high-level cross-modality relations and simultaneously extract low-level detail features as CAI’s modality-specific complementary information. Experiments demonstrate that our method obtains competitive results compared with state-of-the-art methods in visible/infrared image fusion and medical image fusion tasks. Moreover, we surpass other fusion methods in terms of subsequent tasks, averagely scoring 8.27% mAP@0.5 higher in object detection and 5.85% mIoU higher in semantic segmentation.
@@ I Introduction
Multi-modal image fusion has attracted widespread attention because it solves the limitation of single-modality imaging representation. For instance, in the realm of infrared and visible fusion [1, 2, 3], infrared imaging excels in highlighting thermal targets, particularly in challenging environmental conditions, albeit at the expense of resolution. Conversely, visible imaging is adept at preserving intricate texture details but is susceptible to variations in illumination conditions. Consequently, the effective integration of these two modal images enables their respective strengths to be leveraged, such as enhanced visual perception and the identification of salient objects. Fused images exhibit improved adaptability to complex scenarios, particularly when deployed in downstream tasks such as object detection [4] and semantic segmentation [5, 6]. Moreover, image fusion finds extensive applications and extends to various computer vision tasks,
including but not limited to medical image segmentation [7, 8, 9] and remote sensing [10].
@@
In recent years, deep learning methods have become increasingly prevalent in the image fusion field. When dealing with multi-modal images, the primary challenge lies in extracting complementary features from diverse imaging sensors and integrating them into a single image. There are three mainstream methods in multi-modal image fusion: generative methods [11], [12], [13], [14], [15], unified frameworks [16], [17], [18], [19], [20] and auto-encoder methods [21], [22], [23], [24], [25], [26]. Some of these methods have achieved satisfactory performance when tackling specific image fusion tasks. However, existing methods still face three challenges: (i) Most methods do not fully consider the relationships between the two modal images during the image fusion training process; (ii) Local features extracted by convolutional neural networks (CNN) are difficult to control and easy to lose in the forward process, especially high-frequency information; and (iii) Most existing methods often overlook the facilitation of downstream tasks by the generated fused images, resulting in a lack of semantic information and salient targets. Consequently, their performance in downstream tasks is often unsatisfactory.
@@
To tackle the aforementioned challenges, we propose a Semantic-Aware and Multi-Guided Network for Infrared-Visible Image Fusion (SMFNet). For challenge (i), we introduce a Graph Reasoning module to model and reason about high-level relations between two modalities and extract low-level detail features. For challenge (ii), a Cross Attention and Invertible Block (CAI) is employed to address information loss during the decomposing process to extract high-frequency features. Additionally, a base feature extraction module equipped with Multi-Dconv Head Transposed Attention (MDTA) and Local-enhanced Feed Forward (LeFF) network is proposed to capture global features with long-range dependencies. For challenge (iii), we devise two novel loss functions for the two training stages. The loss function in the first stage guides the network to generate an image that closely resembles the two source images. The second loss function, utilized in the second stage, emphasizes the preservation of target intensity and semantic textures in the fused images, with the aim of benefiting downstream tasks. Our motivations, workflow of SMFNet and quantitative results on two datasets are illustrated in Fig. 1. In this paper, our contributions are as follows:
@@
  • A novel loss function based on the Gram matrix is introduced to better utilize semantic details of intensity targets to reconstruct source images. Correlation-driven decomposition loss is proposed to trade off modality-specific, modality-shared and graph-interaction features.
@@
The remainder of this paper is organized as follows. Related works and technologies are described in Section II. Section III demonstrates the semantic-aware and multi-guided network (SMFNet), including encoder, decoder, fusion strategy and loss function. Details about the experiments and evaluation are in Section IV. And Section V is the brief summary of the work.
@@ II-A Infrared-Visible Image Fusion
Traditional approaches have been explored for infrared-visible image fusion in recent years. Multi-scale transform [27], [28], sparse representation-based [29], edge-preserving filter [30], and saliency-based methods [31] are widely recognized as representative methods. Besides, Zhao et al. [32] casted image fusion into a regression problem, seeking the solution in the hierarchical Bayesian manner inferred by expectation-maximization (EM) algorithm. These works have displayed stable performance and have been put into real-world applications. Unavoidably, traditional methods need hand-craft fusion rules and lead to generate artifacts surrounding the edge regions and halos surrounding the objects. The above issues have restricted further development faced with numerous application scenarios.
With the progress of deep-learning technologies, researchers focus on generative methods, unified frameworks and auto-encoder methods in order to explore a series of innovative ways to generate fused images for human vision and machine decision. In generative methods, Ma et al. [12] trained the generator to generate fused images and the discriminator to force fused images similar to the source images. The denoising diffusion probabilistic model (DDPM) [33] was introduced into image fusion with the aim of alleviating the unstable training and lack of interpretability for GAN-based methods. In unified frameworks, Xu et al. [16] proposed the incorporation of elastic weight consolidation into the loss function to enable the model to handle multi-fusion tasks effectively. Zhang et al. [18] offered an alternative approach by uniformly modeling image fusion tasks as the problem of maintaining texture and intensity proportions across source images. Evolved from traditional transforms [34], [35], in auto-encoder methods, Zhao et al. [25, 26] emphasized the encoder’s role in decomposing the image into detailed features and background, with the decoder recovering the original image. Liang et al. [36] devised a pretext task in a self-supervised manner based on common and unique decomposition (CUD). Furthermore, in cooperative training methods, ReCoNet [37] simultaneously learns networks for fusion and registration, aiming to produce robust images for misaligned source images. Liu et al. [38] integrated differentiable neural architecture search (NAS) [39] with bi-level optimization into the training process to suit specific scenario applications. Zhao et al. [40] presents meta-learning as another avenue for joint training, generating features based on the designed network’s capabilities. In MURF [41], image registration and fusion mutually reinforce each other, contributing to enhanced fusion outcomes. Recently, Zheng et al. [42] proposed a new paradigm based on synergistic high-order interaction in both channel and spatial dimensions.
@@ II-B Transformer in Image Processing
In 2020, Dosovitskiy et al. [43] proposed the Vision Transformer (ViT), which represents images as sequences of tokens, similar to how words are represented in natural language processing tasks. It has sparked significant interest and research in the application of Transformer-based architectures to computer vision (CV). Recently, subsequent works [44, 45, 46, 47, 48, 49, 50, 51] have been proposed and applied in various CV tasks, such as image classification, high-resolution image restoration, image super-resolution, deraining and denoising. SegFormer [5] comprising a hierarchical Transformer encoder without positional encoding and lightweight MLP decoders, achieved efficient segmentation on Transformers. SwinIR [44], a strong baseline for image restoration, was equipped with several residual Swin Transformer blocks together with the residual connection as the deep feature extraction module. Wang et al. [45] proposed a non-overlapping window-based self-attention and added a depth-wise convolution layer into the feed-forward network to capture local context. Considering the translation invariance and the vital role of local relationships, Xiao et al. [46] proposed the stochastic window strategy to replace fixed local window strategy, which enjoys powerful representation. XCiT [47] proposed cross-covariance attention, a transposed version of self-attention mechanism, what’s more, the interactions between keys and queries are based on cross-covariance matrix. Restormer[51] focusing on more informative features reduced the time and memory of key-query dot-product interaction in self-attention layers.
@@ II-C Graph Neural Networks
As a typical deep-learning network, Graph Neural Networks (GNN) [52] can model complex links among interpersonal relationships in social networks , thus it has been widely applied in the realm of recommender systems [53]. The Graph Convolutional Networks (GCN) [54] were first proposed for semi-supervised learning on graph-structured data. As for computer vision, Wang et al. [55] answered the question: how do humans recognize the action ”opening a book”, with the definition of graph nodes by the object region from different frames in the video. Exploring the higher performance of semantic segmentation, object detection and instance segmentation tasks, Li et al. [56] proposed to transform a two-dimensional image into a graph structure. The graph’s vertices define clusters of pixels and the edges define the similarity between these regions in a feature space. Han et al. [57] proposed Vision GNN (ViG), viewing image patches as nodes, which introduced graph convolution for aggregating and updating graph information and FFN layers for node feature transformation. To go a step further, Vision HGNN [58] proposed that an image is more than a graph of nodes. They transcended conventional linkages and introduced the concept of hypergraph to encapsulate image information, which is a more universal structure. Recently, Xu et al. [59] proposed a dual-space graph-based interaction network for real-time scene comprehension, which exploited cross-graph and inner-graph relations and extracted contextual information. In RGB-D salient object detection [60], [61], researchers aim at addressing a major technical challenge: how to fully leverage cross-modal complementary features. Correspondingly, GNNs have shown their abilities to learn high-level relations and describe low-level details simultaneously[62].
@@ III-A Overview
As illustrated in Fig. 2, the proposed SMFNet is constructed using an encoder-decoder architecture. To comprehensively extract features, we have meticulously designed an encoder with three branches. Through the encoder, we extract shallow features and subsequently obtain various types of deep features, including local high-frequency information and background low-frequency information. There are three main parallel modules: Cross Attention and Invertible (CAI) block, Base Feature Extraction (BFE) module, and Graph Reasoning (GR) module. Following the aggregation of modality-specific features and modality-shared features from three branches, the decoder reconstructs the original image.
Overall, in training stage I, the encoder’s capability to extract features from the source image and the decoder’s ability to obtain the reconstructed image are trained in an iterative procedure via pursuing minimal reconstruction loss; In training stage II, fusion layers, consistent with the encoding branches, are proposed as the fusion strategy. Constraining the similarity between the source images and the fused image in terms of structure, texture details and intensity, SMFNet can generate an image that benefits visual and computer perception; In the inference phase, the network structure utilized during training stage II can be directly applied to validate the efficacy of image fusion and assess the model’s generalization capability in practical applications.
@@ III-B4 Graph Reasoning (GR) Module
Figure 4: Proposed detail fusion layers in training stage II.
@@ III-D Decoder
The transformer block within the decoder adopts an identical structure to those in the $SFE$ and the base fusion layer. It utilizes depthwise convolution to reconstruct aggregated features from the space $\mathbb{R}^{B\times C\times H\times W}$ to $\mathbb{R}^{H\times W\times 1}$.
@@ III-E Loss Function
$\mathcal{L}_{total}^{II}=\mathcal{L}_{intensity}^{II}+\alpha_{2}\mathcal{L}_{decomp}^{II}+\alpha_{3}\mathcal{L}_{grad}^{II},$ (28)
@@ IV-A Experiment Configuration
As for hyperparameters in Eqs. (23), (24), (25) and (28), $\alpha_{1}$ was uniformly set to 5; $\beta_{1}$ and $\beta_{2}$ were uniformly set to 8 and 10 in training stage I; $\alpha_{2}$ and $\alpha_{3}$ were set to 10 and 2 in training stage II. We used the Adam optimizer, and the initial learning rate was ${1e}^{-4}$ without weight decay. Experiments were executed on a PC with two NVIDIA GeForce 4090 GPUs and the pytorch framework.
@@ IV-C2 Quantitative comparison
In Tab. I, we present eight evaluation metrics across three test sets for quantitative analysis. Specifically, the highest EN and AG indicate significant preservation of information from the source images in our fused images, thanks to the cross-attention invertible module and effective complementary feature aggregation. The attention to targeted edge contours is largely attributed to the semantic loss function. Additionally, the highest values of VIF and Qabf indicate that the generated fused images exhibit excellent visual performance suitable for the human visual system and are also well-suited for subsequent computer vision tasks. Furthermore, the highest SF demonstrates good texture details in our results.
TABLE I: Quantitative evaluation results of IVF task, where red indicates the best results, and blue shows the second-best values.

### v3

@@ Abstract
Multi-modality image fusion aims at fusing modality-specific (complementarity) and modality-shared (correlation) information from multiple source images. To tackle the problem of the neglect of inter-feature relationships, high-frequency information loss, and the limited attention to downstream tasks, this paper focuses on how to model correlation-driven decomposing features and reason high-level graph representation by efficiently extracting complementary information and aggregating multi-guided features. We propose a three-branch encoder-decoder architecture along with corresponding fusion layers as the fusion strategy. Firstly, shallow features from individual modalities are extracted by a depthwise convolution layer combined with the transformer block. In the three parallel branches of the encoder, Cross Attention and Invertible Block (CAI) extracts local features and preserves high-frequency texture details. Base Feature Extraction Module (BFE) captures long-range dependencies and enhances modality-shared information. Graph Reasoning Module (GR) is introduced to reason high-level cross-modality relations and simultaneously extract low-level detail features as CAI’s modality-specific complementary information. Experiments demonstrate the competitive results compared with state-of-the-art methods in visible/infrared image fusion and medical image fusion tasks. Moreover, the proposed algorithm surpasses the state-of-the-art methods in terms of subsequent tasks, averagely scoring 8.27% mAP@0.5 higher in object detection and 5.85% mIoU higher in semantic segmentation. The code is avaliable at https://github.com/Abraham-Einstein/SMFNet/.
@@ I Introduction
Multi-modal image fusion has attracted widespread attention because it solves the limitation of single-modality imaging representation. For instance, in the realm of infrared and visible fusion [1, 2, 3], infrared imaging excels in highlighting thermal targets, particularly in challenging environmental conditions, albeit at the expense of resolution. Conversely, visible imaging is adept at preserving intricate texture details but is susceptible to variations in illumination conditions. Consequently, the effective integration of these two imaging modalities allows their respective strengths to be fully utilized, such as enhancing visual perception and improving the identification of salient objects. Fused images exhibit improved adaptability to complex scenarios, particularly when deployed in downstream tasks such as object detection [4] and semantic segmentation [5, 6]. Moreover, image fusion finds extensive applications and extends to various computer vision tasks, including but not limited to medical image segmentation [7, 8, 9], remote sensing [10].
@@
In recent years, deep learning methods have become increasingly prevalent in the image fusion field. The primary challenge of image fusion lies in extracting complementary features from diverse imaging sensors and integrating them into a single image. There are three mainstream methods in multi-modal image fusion: generative methods [11], [12], [13] [14], unified frameworks [15], [16], [17], [18] and auto-encoder methods [19], [20], [21], [22]. It has been verified that the effective embedding strategy preserves details and enhances the overall image style [23]. Some multi-branch guided methods [24],[25] have also achieved satisfactory performance when tackling specific image fusion tasks. However, existing methods still face three challenges: (i) Most methods do not fully consider the relationships between the two modal images during the image fusion training process; (ii) Local features extracted by convolutional neural networks (CNN) are easy to lose in the forward process, especially high-frequency information; and (iii) Most existing methods often overlook the facilitation of downstream tasks by the generated fused images, resulting in a lack of semantic information and salient targets. Consequently, their performance in downstream tasks is often unsatisfactory.
@@
To tackle the aforementioned challenges, we propose a Semantic-Aware and Multi-Guided Network for Infrared-Visible Image Fusion (SMFNet). For challenge (i), a Graph Reasoning module is introduced to model and reason high-level relations between two modalities and extract low-level detail features. For challenge (ii), a Cross Attention and Invertible Block (CAI) is employed to extract high-frequency features. And a base feature extraction module equipped with Multi-Dconv Head Transposed Attention (MDTA) and Local-enhanced Feed Forward (LeFF) network is proposed to capture global features with long-range dependencies. For challenge (iii), we devise two novel loss functions for the two training stages. The first stage loss function guides the network to reconstruct source images. The second stage loss function emphasizes the preservation of target intensity and semantic textures in the fused images, intending to benefit downstream tasks. The workflow of AE-based methods, SMFNet, and quantitative results on two datasets are illustrated in Fig. 1. In this paper, our contributions are as follows:
@@
  • A novel loss function based on the Gram matrix is introduced to better utilize the semantic details of intensity targets to reconstruct source images. Correlation-driven decomposition loss is proposed to trade off modality-specific, modality-shared and graph-interaction features.
@@
The remainder of this paper is organized as follows. Related works and technologies are described in Section II. Section III demonstrates the semantic-aware and multi-guided network (SMFNet). Details about the experiments and evaluation are in Section IV. And Section V is the brief summary of the work.
@@ II-A Infrared-Visible Image Fusion
Traditional approaches have been explored for infrared-visible image fusion in recent years. Multi-scale transform [26], [27], sparse representation-based [28], edge-preserving filter [29], and saliency-based methods [30] are widely recognized as representative methods. Unavoidably, traditional methods need hand-crafted fusion rules, which have restricted further development faced with numerous application scenarios.
With the progress of deep-learning technologies, researchers focus on generative methods, unified frameworks and auto-encoder methods to generate fused images for human vision and machine decision. In generative methods, Ma et al. [12] trained the generator to generate fused images and the discriminator to force fused images similar to the source images. The denoising diffusion probabilistic model (DDPM) [31] was introduced into image fusion to alleviate the unstable training for GAN-based methods. One of the representative diffusion-based works is VDMUFusion [32], which integrated the fusion problem into the forward and reverse processes allowing for predicting the noise and fusion mask simultaneously. In unified frameworks, Xu et al. [15] proposed the incorporation of elastic weight consolidation into the loss function to enable the model to handle multi-fusion tasks effectively. Zhang et al. [17] offered an alternative approach by uniformly modeling image fusion tasks as the problem of maintaining texture and intensity proportions across source images. Evolved from traditional transforms [33], [34], auto-encoder methods [21, 22, 35] emphasized the encoder’s role in decomposing the image into detailed features and background, with the decoder recovering the original image. Furthermore, in cooperative training methods, ReCoNet [36] simultaneously learns networks for fusion and registration, aiming to produce robust images for misaligned source images. In MURF [37], image registration and fusion mutually reinforce each other, contributing to enhanced fusion outcomes. As an end-to-end training scheme, Yao et al. [23] proposed a novel fusion network with hierarchical guidance, which obtained salient maps through Laplacian Pyramid decomposition and achieved pixel-level saliency preservation. Recently, Zheng et al. [38] proposed a new paradigm based on synergistic high-order interaction in both channel and spatial dimensions. Yao et al. [24] combined the uncertainty parameters from the segmentation network to achieve enhancement and fusion process. DLN-CFN [25], considering the content and color information, fused visible and infrared images in the low-light scenarios.
@@ II-B Transformer in Image Processing
Considering the translation and rotation invariance of image properties, transformer was introduced from the field of NLP to image processing, seeking unified modeling of natural language and images. In 2020, Dosovitskiy et al. [39] proposed the Vision Transformer (ViT), which represents images as sequences of tokens, similar to how words are represented in natural language processing tasks. It has sparked significant interest and research in the application of Transformer-based architectures to computer vision (CV). Recently, subsequent works [40, 44, 45, 46, 47] have been proposed and applied in various CV tasks, such as image classification, high-resolution image restoration, image super-resolution, deraining and denoising. SegFormer [5] comprising a hierarchical Transformer encoder without positional encoding and lightweight MLP decoders, achieved efficient segmentation on Transformers. SwinIR [40], a strong baseline for image restoration, was equipped with several residual Swin Transformer blocks together with the residual connection as the deep feature extraction module. Wang et al. [41] proposed a non-overlapping window-based self-attention and added a depth-wise convolution layer into the feed-forward network to capture local context. Considering the translation invariance and the vital role of local relationships, Xiao et al. [42] proposed the stochastic window strategy to replace fixed local window strategy, which enjoys powerful representation. XCiT [43] proposed cross-covariance attention, a transposed version of self-attention mechanism, what’s more, the interactions between keys and queries are based on cross-covariance matrix. Restormer[47] focusing on more informative features reduced the time and memory of key-query dot-product interaction in self-attention layers. Variants of transformers also has been applied for medical and biological image processing. For medical image fusion, MM-Net [48] highlighted the strengths of the MixFormer [49] backbone and also fully utilized the multi-scale local features and context information.
@@ II-C Graph Neural Networks
As a typical deep-learning network, Graph Neural Networks (GNN) [50] can model complex links among interpersonal relationships in social networks The Graph Convolutional Networks (GCN) [52] were first proposed for semi-supervised learning on graph-structured data. As for computer vision, Wang et al. [53] answered the question: how do humans recognize the action ”opening a book”, with the definition of graph nodes by the object region from different frames in the video. Exploring the higher performance of semantic segmentation, object detection and instance segmentation tasks, Li et al. [54] proposed to transform a two-dimensional image into a graph structure. The graph’s vertices define clusters of pixels and the edges define the similarity between these regions in a feature space. Han et al. [55] proposed Vision GNN (ViG), viewing image patches as nodes, which introduced graph convolution for aggregating and updating graph information and FFN layers for node feature transformation. To go a step further, Vision HGNN [56] proposed that an image is more than a graph of nodes. They transcended conventional linkages and introduced the concept of hypergraph to encapsulate image information, which is a more universal structure. Recently, Xu et al. [57] proposed a dual-space graph-based interaction network for real-time scene comprehension, which exploited cross-graph and inner-graph relations and extracted contextual information. In RGB-D salient object detection [58], [59], researchers aim at addressing a major technical challenge: how to fully leverage cross-modal complementary features. Correspondingly, GNNs have shown their abilities to learn high-level relations and describe low-level details simultaneously[60].
@@ III-A Overview
As illustrated in Fig. 2, the proposed SMFNet is constructed using an encoder-decoder architecture. Overall, in training stage I, the encoder’s capability to extract features from the source image and the decoder’s ability to obtain the reconstructed image are trained in an iterative procedure via pursuing minimal reconstruction loss; In training stage II, fusion layers, consistent with the encoding branches, are proposed as the fusion strategy. The identical structure of the two streams shares parameters throughout the entire training process. Constraining the similarity between the source images and the fused image in terms of structure, texture details and intensity, SMFNet can generate an image that benefits visual and computer perception; In the inference phase, the network structure utilized during training stage II can be directly applied to validate the efficacy of image fusion and assess the model’s generalization capability in practical applications.
@@ III-B3 Base Feature Extraction (BFE) Module
Figure 4: Proposed detail fusion layers in training stage II.
@@ III-D Decoder
The transformer block within the decoder adopts an identical structure to those in the $SFE$ and the base fusion layer. It utilizes depthwise convolution to reconstruct aggregated features from the space $\mathbb{R}^{B\times C\times H\times W}$ to $\mathbb{R}^{H\times W\times 1}$. Then the fused image $\mathbb{R}^{H\times W\times 3}$ is obtained via YCbCr space conversion [22]. The details of the whole network are listed in Tab. I.
TABLE I: Details architecture of our proposed SMFNet. Head refers to multi-head self-attention in a Transformer model. The terms I, O, and K correspond to the input dimension, output dimension, and kernel size, respectively.
Layer Explanation Module
Encoder Fig. 2(a) DConv(I1O64K3)
Transformer(I64O64Head8)
Fig. 2(c) CAI Block(I64O64)
Fig. 2(d) BFE Block(I64O64)
Fig. 2(e) GR Module(I64O64)
$F_{detail}$ Fig. 4 Detail Module(I64O64)
$F_{base}$ Transformer(I64O64Head8)
$F_{graph}$ Conv(I128O64K1)
Decoder Fig. 2(a) Transformer(I64O64Head8)
DConv(I64O64K3)
DConv(I64O64K3)
DConv(I64O64K3), Sigmoid
@@ III-E Loss Function
$\mathcal{L}_{total}^{II}=\mathcal{L}_{intensity}^{II}+\alpha_{2}\mathcal{L}_{grad}^{II}+\alpha_{3}\mathcal{L}_{decomp}^{II},$ (28)
@@ IV-A Experiment Configuration
As for hyperparameters in Eqs. (23), (24), (25) and (28), $\alpha_{1}$ was uniformly set to 2; $\beta_{1}$ and $\beta_{2}$ were uniformly set to 8 and 10 in training stage I; $\alpha_{2}$ and $\alpha_{3}$ were set to 10 and 2 in training stage II. We used the Adam optimizer, and the initial learning rate was ${1e}^{-4}$ without weight decay. Experiments were executed on a PC with two NVIDIA GeForce 4090 GPUs and the pytorch framework.
@@ IV-C2 Quantitative comparison
In Tab. II, we present eight evaluation metrics across three test sets for quantitative analysis. Specifically, the highest EN and AG indicate significant preservation of information from the source images in our fused images, thanks to the cross-attention invertible module and effective complementary feature aggregation. The attention to targeted edge contours is largely attributed to the semantic loss function. Additionally, the highest values of VIF and Qabf indicate that the generated fused images exhibit excellent visual performance suitable for the human visual system and are also well-suited for subsequent computer vision tasks. Furthermore, the highest SF demonstrates good texture details in our results.
TABLE II: Quantitative evaluation results of IVF task, where red indicates the best results, and blue shows the second-best values.